Accelerator Update for March 2, 2009 - March 4, 2009 |
Monday March 2 The day shift began with Operations monitoring the Tevatron (TeV) store 6850 (89.69E30) and the Recycler stash (259.9E10), with the Antiproton Source stacking (14.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. At 9:36 AM, a Booster RF station (BRF15) tripped off for the first of two times in 34 minutes. It reset. At 10:16 AM, BRF4 and 9 tripped off. They reset. At 10:37 AM, BRF8 tripped off. It reset. At 10:50 AM, two Booster vacuum valves (VB13 and VB17) closed. At 10:57 AM, the Booster Gradient Magnet Power Supply tripped off. At 11:03 AM, Operations received a West Gallery LCW low pressure alarm. The Booster LCW pump at CUB had tripped off. Operations contacted the Duty Mechanic. Operations put Pbar into standby at 11:09 AM. TeV experts began a collimator end-of-store study at 11:11 AM. Operations reported at 11:16 AM that the Booster LCW system had resumed normal operating pressures. At 11:53 AM, the 400 MeV area beam permit tripped off due to bad beam positions. It reset. Operations terminated store 6850 at 12:32 PM. Operations put the TeV into shot setup at 12:48 PM. At 12:56 PM, the 400 MeV area tripped off due to bad beam positions. It reset. Operations established store 6851 at 2:56 PM with an initial luminosity of 310.32E30. At 3:59 PM, the 400 MeV area tripped off. It reset. Operations reported at 4:08 PM that the Recycler had stashed antiprotons six times during the day shift. At 4:18 PM, the 400 MeV area tripped off for the first of four times in 33 minutes. An operator was tuning the Linac during these trips. At 5:22 PM, a Pbar Debuncher power supply (D:H926) tripped off. Operations contacted and expert who said the trip was due to a PLC failure. Operations called in an expert. A power supply expert reported at 6:57 PM that he had loaded new software for the D:H926 PLC. At 7:44 PM, LRF3 tripped off. It reset. Operations reported at midnight that the Recycler had stashed six times during the evening shift. Tuesday March 3 The midnight shift began with Operations monitoring store 6851 (105.97E30) and the stash (252.0E10), with Pbar stacking (26.2E10), and with NuMI and MiniBooNE taking beam. At 12:30 AM, LRF3 tripped off. It reset. At 12:53 AM, a Linac 750 MeV area quadrupole power supply (L:HQTM4U) had gone to zero and the H- Source tripped off. Operators tried to turn the power supply on, but there was no remote control. Operations put Pbar into standby at 1:10 AM. D0 reported at 1:11 AM that they were conducting luminosity studies. Operations reported at 1:27 AM that they had turned on L:HQTM4U locally. At 1:54 AM, HQTM4U tripped off again. It reset. Operations called in an expert. At 2:12 AM, HQTM4U tripped off. It reset. The H- Source also tripped off. It reset. Operations terminated store 6851 at 2:38 AM. Operations prepared D0 and CDF for access at 2:53 AM. CDF came out of access at 3:30 AM, D0 came out of access at 3:51 AM. Operations put the TeV through a dry squeeze at 4:27 AM. Operations reported at 4:55 AM that a power supply expert had replaced the HQTM4U power supply. Operations put the TeV into shot setup at 6:12 AM. At 6:35 AM, a Booster vacuum pump (IPRF17) tripped off and caused BRF7 and 8 to trip off. Everything reset. At 6:44 AM, BRF19 tripped off. It reset. At 7 AM, BRF5 tripped off. It reset. At 7:06 AM, BRF15 tripped off. It reset. At 7:09 AM, BRF15 and 16 tripped off. It reset. Operations established store 6855 at 7:54 AM with an initial luminosity of 310.98E30. Operations reported at 8:05 AM that the Recycler had stashed five times during the midnight shift. At 8:31 AM, the 400 MeV area tripped off for the first of seven times in 37 minutes. It reset. An expert tuned. At 11:39 AM, the Neutron Therapy Facility (NTF) reported that their beam had stopped. Experts investigated and found a CAMAC crate with a power supply failure. Linac experts reported at 1:01 PM that they had replaced the failed crate power supply. NTF resumed treating patients. At 2:39 PM, the NuMI beam permit tripped off. It reset. At 4:08 PM, Operations reported that an Accumulator kicker (A:IKIKP1) stopped working and would not reset. Operators investigated and found its breaker had tripped. Operations reported at 4:36 PM that the Recycler had stashed seven times during the midnight shift. Operations terminated store 6855 at 10:20 PM. Operations put the TeV into shot setup at 10:32 PM. At 11:56 PM, the TeV suffered a quench at sectors C4, DA, DB, and DD. Operations investigated. Operations reported at midnight that the Recycler had stashed seven times during the evening shift. Wednesday March 4 The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash (8.5E10), with Pbar stacking (59.6E10), and with NuMI and MiniBooNE taking beam. At 12:40 AM, Operations reported a MI power supply (HTF11) trip, which was probably caused by the auto tune program. At 2:20 AM, the P1 transfer line tripped off. It reset. Operations put the TeV through a dry squeeze at 4:29 AM. Operations reported at 5:02 AM that TeV experts were investigating the quench. CUB experts reported at 7:43 AM that they had taken column 1 off line for regeneration. Operations reported at 7:53 AM that the Recycler had stashed seven times during the midnight shift. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 03/04/2009 email Fermilab |
| Security, Privacy, Legal | |